Statistics department data show that 20% of Pakistani age 18 and over holds a certificate of matriculation. If 10 Pakistani national age 18 and above are randomly selected, find the probability that at least 3 will have certificates of matriculation.
1
Expert's answer
2020-06-15T19:12:26-0400

Given :

n=10

p=0.2


conditions for binomial distribution:

The number of observations n is fixed. 2: Each observation is independent. 3: Each observation represents one of two outcomes ("success" or "failure"). 4: The probability of "success" p is the same for each outcome

since above conditions are satisfied , we use binomial distribution.


To find : 

"P(x\geq 3)"


Using binomial table for n= 10, p=0.2,x=3,



"P(x\geq 3)"  = 1- P( x<3)= 1- 0.6778= 0.3222